Solve (D2 + a2)y = sinax
1
Expert's answer
2021-07-27T17:12:04-0400

Let us solve the differential equation "(D^2 + a^2)y = \sin ax." The characteristic equation "k^2+a^2=0" has the solutions "k_1=ai, k_2=-ai," and hence the general solution of the homogeneous differential equation is "y=C_1\cos ax+ C_2\sin ax." The particular solution of non-homogeneous equation is of the form "y_p=x(L\sin ax+ M\cos ax)." Then "y_p'=L\sin ax+ M\cos ax+x(aL\cos ax-a M\sin ax)," and "y_p''=aL\cos ax-a M\sin ax+aL\cos ax-a M\sin ax+x(-a^2L\sin ax-a^2 M\cos ax)."

It follows that "2aL\cos ax-2a M\sin ax+x(-a^2L\sin ax-a^2 M\cos ax)+a^2x(L\sin ax+ M\cos ax)=\sin ax,"

and hence "2aL=0" and "-2aM=1." Therefore, if "a\ne 0" then "L=0,M=-\frac{1}{2a}."


We conclude that if if "a\ne 0" then the general solution of the differential equation "(D^2 + a^2)y = \sin ax" is "y=C_1\cos ax+ C_2\sin ax-\frac{1}{2a}x\cos ax." If "a=0" then the general solution of the differential equation "D^2 y = 0" is "y=C_1x+C_2."